systems of linear equations and inequalities, please help me solve these equations, i'm stumped
#1 v-w=5
v+2w=4
#2 2g-3h=0
3g-2h=5

Thank you, Jack

Answer by mananth(16946)

v-w=5..........................1
v+2w=4.........................2
..
multiply equation 1 by -1
-v+w=-5
add to (2)
3w=-1
w= -1/3
...
plug value of w in (1)
v+1/3 = 5
v= 5-1/3
v=14/3
...
2g-3h=0...................1
3g-2h=5....................2
..
multiply (1) by 3 & (2) by -2
6g-9h=0
-6g+4h=-10
Add them up
-5h=-10
h=2
..
plug h in (1)
2g-3h=0
2g-6=0
2g=6
g= 3
..
m.ananth@hotmail.ca